How do I solve
7/x = 9/x - 1/11?
6/z+8 = 5/z-3?

4-s/7-s = 3/s-7?

Answer by stanbon(75887)

How do I solve
7/x = 9/x - 1/11?
Multiply thru by 11x to get:
77 = 99 - x
x = 99-77
x = 22
---------------------

6/z+8 = 5/z-3?
Cross-multiply:
6(z-3) = 5(z+8)
6z-18 = 5z+40
z = 58
--------------------------
4-s/7-s = 3/s-7?
--
Modify the fraction on the left by multiplying numerator and denominator by -1.
{(s-4)/(s-7)]= 3/(s-7)
Multiply both sides by s-7 to get:
s-4 = 3
s = 7
But s cannot be seven because the denominator is s-7 and
that would be zero if s = 7.
Conclusion: No solution
=================================
Cheers,
Stan H.